[figure: 9.7 cm]

REVERSIBLE BEHAVIOR OF MAGNETOELASTIC DOMAIN STRUCTURE IN THE LAYERED ANTIFERROMAGNET OF

CoCl_{2}-TYPE

E.V. Gomonaj^{a}, V. M. Loktev^{a}, S.M. Ryabchenko^{b}, A.F. Lozenko^{b}, P.A.

Trotsenko^{b}, V.M. Kalita^{b} aNational Technical University of Ukraine ”KPI”, 37, ave Peremogy Kyiv, 252056, Ukraine bInstitute for Physics, NAS of Ukraine, 41, ave Nauki Kyev, 252056, Ukraine

Macroscopic magnetoelastic domain structure of the defectless layered antiferromagnet of CoCl_{2}-type with the ”easy-plane” magnetic anisotropy is studied theoretically and experimentally. Magnetostriction behavior in the external magnetic field applied perpendicular to the main symmetry axis of the crystal points to the reversible behavior of the domain structure. In assumption of mobile domain walls, the finite-size effects are shown to result in formation of stable domain structure that changes reversibly under the action of the external magnetic field and can be treated as equilibrium. It is found that the domain structure and its collective behavior are governed by elasticity. Calculated field dependence of a domain structure, magnetostriction and low-frequency AFMR of poly- and monodomain sample is in qualitative agreement with the available experimental data.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : Helen Gomonaj

Address for correspondence : 18, Klovskii spusk, apt.74 252021 Kyiv, Ukraine Phone: +38(044) 241-77-12 Fax: +38(044) 241-77-12

Email address : malyshen@ukrpack.net